Thirty-one small communities in Connecticut (population near 10,000 each) gave an average of x = 138.5 reported cases of larceny per year. Assume that σ is known to be 43.5cases per year.

(a) Find a 90% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)

(b) Find a 95% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)

(c) Find a 99% confidence interval for the population mean annual number of reported larceny cases in such communities. What is the margin of error? (Round your answers to one decimal place.)
Answer #1

Solution :

Given that,

(a)

Sample size = n = 31

Z\alpha/2 = 1.645

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.645 * (43.5 / \sqrt 31)

Margin of error = E = 12.9

(b)

Sample size = n = 31

Z\alpha/2 = 1.96

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.96 * (43.5 / \sqrt 31)

Margin of error = E = 15.3

(c)

Sample size = n = 31

Z\alpha/2 = 2.576

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 2.576 * (43.5 / \sqrt 31)

Margin of error = E = 20.1